PLEASE ANSWER NUMBER THREE (not number two)

Determine values of the cumulative distribution function for the random variable in the previous problem. 3.

This problem (number 2) is only for reference.

2. The probability mass function below is defined for x 0, 1,2,3,.. fr 5 5 -56 What is the probability for each of the following expressions? a) P(X 2) b) P(XE 2) c) P(X> 2) d) P(X2 1)
Answer #1

3)here as cumulative distribution function:

F(x)=P(X<=x)=P(X=0)+P(X=1)+P(X=2)+...+P(X=x)

=(3/5)*((2/5)0+(2/5)1+(2/5)2+...(2/5)x)

=(3/5)*1*(1-(2/5)x+1)/(1-2/5)

F(x)=1-(2/5)x+1 for x=0,1,2,3,.............